PROBLEM 3 Consider equally risky, all-equity financed firms G and D. Both firms are currently trading at $50 per share. Firm G pays no dividends, but firm D pays all its earnings as dividends. In a year from today, the stock of firm G is expected to be at $65 per share. Firm D is expected to pay a $15 dividend per share at the year-end and its stock price at the year-end is expected to be $50 per share. Capital gains are taxed at half the 40% regular income tax rate, but the dividend income is taxed at 30%.

Required: At what share price should the stock of firm D be trading to-day in a competitive stock market? Show calculations to support your answer.
Answer #1

Since the risk return profile of both the cos are same, the Return provided by both the Firms will be same. They can be compared. The only difference is that one co gives out Dividend but other gives the Capital Gain after retaining all the Dividend and reinvesting it.

Firm D Dividend after tax = 15 (1-0.30) = 10.5

Firm G Capital Gain after tax = 15 * (1-0.40/2) = 12

Return given by the Firm D = After tax Capital Gain / Price

= 12 / 50

= 24%

Now, Since the firms are simimar, we can equate the return from one firm to other. i.e. The return of Firm D can be equated with the Return of Firm G to calculate the Price of Firm G.

Therefore,

Return of Firm G = After tax Dividend / Price

24% = 10.5 / Price

Price = 10.5 / 0.24

Price =$43.75